A 500 kg elevator starts from rest. It moves upward for 2.50 s with constant acceleration until it reaches its cruising speed, 1.75 m/s.

a.) What is the average power of the elevator motor during this period?

b.) How does this power compare with the motor power when the elevator moves at its cruising speed?
Answer #1

given

m = 500 kg

u = 0

v = 1.75 m/s

t = 2.5 s


let a is the acceleration.

apply, v = u + a*t

a = (v - u)/t

= (1.75 - 0)/2.5

= 0.7 m/s^2

Force exerted by elevator engine = m*g + m*a

= 500*9.8 + 500*0.7

= 5250 N

distance travelled during this time, d = u*t + 0.5*a*t^2

= 0 + 0.5*0.7*2.5^2

= 2.1875

a) the average power of the elevator motor during this period,

P_avg(during starting) = workdone/time taken

= F*d/t

= 5250*2.1875/2.5

= 4594 Watts

b) Pavg(at crusing speed) = F*v

= 5250*1.75

= 9187.5 watts

P_avg(during starting)/Pavg(at crusing speed) = 4594/9187.5


P_avg(during starting) = 0.5* Pavg(at crusing speed)